OpenStudy (anonymous):

solve for y. Z=5/6(x+5y)

OpenStudy (anonymous):

Multiply by x+5y \[z(x+5y)=\frac{5}{6}\] Distribute the z \[zx+5zy=\frac{5}{6}\] Isolate the y term \[5zy=\frac{5}{6}-zx\] Divide by 5z \[y=\frac{1}{5z}(\frac{5}{6}-zx)\] Distribute \[y=\frac{1}{6z}-\frac{x}{5}\]